# The ratio monotonicity of Eulerian-type polynomials

### v1

@@ 1 Introduction
Let $a(x)=\sum_{i=0}^{n}a_{i}x^{i}$ be a polynomial with nonnegative coefficients. We say that $a(x)$ is unimodal if $a_{0}\leqslant a_{1}\leqslant\cdots\leqslant a_{k-1}\leqslant a_{k}\geqslant a_{k+1}\geqslant\cdots\geqslant a_{n}$ for some $k$, where the index $k$ is called the mode of $a(x)$. It is log-concave if $a_{i}^{2}\geqslant a_{i-1}a_{i+1}$ for all $1\leqslant i\leqslant n-1$. Clearly, a log-concave sequence with no internal zeros is unimodal, but the converse is not true. The polynomial $a(x)$ is said to be gamma-positive if there exist nonnegative numbers $\gamma_{k}$ such that
@@
We say that $a(x)$ is spiral if $a_{n}\leqslant a_{0}\leqslant a_{n-1}\leqslant a_{1}\leqslant\cdots\leqslant a_{\lfloor n/2\rfloor}$. Following [26, Definition 2.9], the polynomial $a(x)$ is alternatingly increasing if $a_{0}\leqslant a_{n}\leqslant a_{1}\leqslant a_{n-1}\leqslant\cdots\leqslant a_{\lfloor{(n+1)}/{2}\rfloor}$. Clearly, if $a(x)$ is spiral and $\deg a(x)=n$, then $x^{n}a(1/x)$ is alternatingly increasing, and vice versa. The notion of ratio monotonicity for polynomials was introduced by Chen-Xia [12]. We say that $a(x)$ is ratio monotone if
@@
Clearly, $a(x)$ and $b(x)$ are both symmetric. Moreover, $\deg a(x)=1+\deg b(x)$ or $b(x)=0$. Following Brändén-Solus [5], we call the ordered pair of polynomials $(a(x),b(x))$ the symmetric decomposition of $f(x)$. According to [21, Definition 1.2], if $a(x)$ and $b(x)$ are both gamma-positive, then $f(x)$ is said to be bi-gamma-positive. Bi-gamma-positivity implies alternatingly increasing property and gamma-positivity can be seen as a degeneration of bi-gamma-positivity. For example, the polynomial $1+10x+4x^{2}$ is bi-gamma-positive, since
@@ 2.1 A proof of the bi-gamma-positivity of $f_{n}(x)$
with $\alpha_{0,0}=1,~\beta_{0,0}=0$ and $\alpha_{0,k}=\beta_{0,k}=0$ for $k\geqslant 1$. Therefore, when $a+c\geqslant b\geqslant c>0$, we see that $f_{n}(x)$ is bi-gamma-positive, and when $a>0$ and $b=c$, $f_{n}(x)$ is gamma-positive. ∎
Suppose that $p(x)$ and $q(x)$ have only real zeros, the zeros of $p(x)$ are $\xi_{1}\leqslant\cdots\leqslant\xi_{n}$, and that those of $q(x)$ are $\theta_{1}\leqslant\cdots\leqslant\theta_{m}$. Following [14], we say that $p(x)$ interlaces $q(x)$ if $\deg q(x)=1+\deg p(x)$ and the zeros of $p(x)$ and $q(x)$ satisfy $\theta_{1}\leqslant\xi_{1}\leqslant\theta_{2}\leqslant\cdots\leqslant\xi_{n}\leqslant\theta_{n+1}$. We use the notation $p(x)\prec q(x)$ for $p(x)$ interlaces $q(x)$. Combining (6) and [18, Theorem 2.3], one can immediately get the following result.
@@
Let $\{f_{n}(x)\}_{n\geqslant 0}$ be a sequence of polynomials satisfying (4). Then $f_{n}(x)$ has only real zeros. Suppose that $(a_{n}(x),b_{n}(x))$ is the symmetric decomposition of $f_{n}(x)$. Then both $a_{n}(x)$ and $b_{n}(x)$ have only real zeros and $b_{n}(x)\prec a_{n}(x)$.
@@ 2.2 A proof of the ratio monotonicity of $f_{n}(x)$
Comparing with (8) and (9), we immediately get the desired result. This completes the proof. ∎
@@ 3.1 $q$-Eulerian polynomials, $1/k$-Eulerian polynomials and generalized Eulerian polynomials
Setting $L_{n}(x)=x^{n}A_{n+1}(1/x,q)$, it follows from (5) and (17) that
$L_{n+1}(x)=(nx+qx+1)L_{n}(x)+x(1-x)\frac{\mathrm{d}}{\mathrm{d}x}L_{n}(x).$ (18)
Following Hwang-Chern-Duh [17, p. 26], the polynomials $L_{n}(x)$ can be called LI Shanlan polynomials, since these polynomials appeared in his 1867 book. Combining the recursions (17), (18) and Theorem 1, we can give the following result.
@@
It should be noted that $A_{n}(x,2)$ is the big descent polynomials over $\mathfrak{S}_{n+1}$, where a big descent is an index $i$ such that $\pi(i)\geqslant\pi(i+1)+2$, see [27, A120434]. We list the first few polynomials:
@@
The bi-gamma-positivity of $A_{n}^{(k)}(x)$ was first established in [21, Section 3.4], and Yan-Yang-Lin [30] gave a nice combinatorial proof of this result. Combining Corollary 4 and (19), we get the following result.
@@
These polynomials were introduced by Morisita [23], and they were also studied by Carlitz-Scoville [9]. They also appeared in the context of random staircase tableaux [16]. Hitczenko-Janson [16] investigated the asymptotic distribution of these polynomials. Combining (20) and Theorem 1, we end this subsection by giving the following result.
@@ 3.3 The $r$-colored Eulerian polynomials
When $r=1$ and $r=2$, the polynomial $A_{n,r}(x)$ reduces to the types $A$ and $B$ Eulerian polynomials $A_{n}(x)$ and $B_{n}(x)$, respectively. Very recently, there has been much work devoted to the bi-gamma-positivity of the $r$-colored Eulerian polynomials and their variations, see [2, 15, 21] for details. In particular, it is now well known that $A_{n,r}(x)$ is bi-gamma-positive when $r>2$, see [2, Eq. (21)] and [21, Theorem 7.5].

### v3

@@ 1 Introduction
Let $a(x)=\sum_{i=0}^{n}a_{i}x^{i}$ be a polynomial with nonnegative coefficients. We say that $a(x)$ is unimodal if $a_{0}\leqslant a_{1}\leqslant\cdots\leqslant a_{k-1}\leqslant a_{k}\geqslant a_{k+1}\geqslant\cdots\geqslant a_{n}$ for some $k$, where the index $k$ is called the mode of $a(x)$. It is log-concave if $a_{i}^{2}\geqslant a_{i-1}a_{i+1}$ for all $1\leqslant i\leqslant n-1$. Clearly, a log-concave sequence with no internal zeros is unimodal, but the converse is not true. Darroch [14] showed that if $a(x)$ has only real nonpositive zeros, then $a(x)$ is unimodal with at most two modes and each mode $k$ satisfies
$\left\lfloor\frac{a^{\prime}(1)}{a(1)}\right\rfloor\leqslant k\leqslant\left\lceil\frac{a^{\prime}(1)}{a(1)}\right\rceil.$
The polynomial $a(x)$ is called gamma-positive if there exist nonnegative numbers $\gamma_{k}$ such that
@@
We say that $a(x)$ is spiral if $a_{n}\leqslant a_{0}\leqslant a_{n-1}\leqslant a_{1}\leqslant\cdots\leqslant a_{\lfloor n/2\rfloor}$. Following [27, Definition 2.9], the polynomial $a(x)$ is alternatingly increasing if $a_{0}\leqslant a_{n}\leqslant a_{1}\leqslant a_{n-1}\leqslant\cdots\leqslant a_{\lfloor{(n+1)}/{2}\rfloor}$. Clearly, if $a(x)$ is spiral and $\deg a(x)=n$, then $x^{n}a(1/x)$ is alternatingly increasing, and vice versa. The notion of ratio monotonicity for polynomials was introduced by Chen-Xia [12]. We say that $a(x)$ is ratio monotone if
@@
Clearly, $a(x)$ and $b(x)$ are both symmetric. Moreover, $\deg a(x)=1+\deg b(x)$ or $b(x)=0$. Following Brändén-Solus [5], we call the ordered pair of polynomials $(a(x),b(x))$ the symmetric decomposition of $f(x)$. According to [22, Definition 1.2], if $a(x)$ and $b(x)$ are both gamma-positive, then $f(x)$ is said to be bi-gamma-positive. Bi-gamma-positivity implies alternatingly increasing property and gamma-positivity can be seen as a degeneration of bi-gamma-positivity. For example, the polynomial $1+10x+4x^{2}$ is bi-gamma-positive, since
@@
As Hwang-Chern-Duh [18, Section 4.1] once said, one of the most common recursions with rich combinatorial properties among the extensions of Eulerian numbers is given as follows:
$\mathcal{P}_{n+1}(x)=(qnx+(qr-p)x+p)\mathcal{P}_{n}(x)+qx(1-x)\frac{\mathrm{d}}{\mathrm{d}x}\mathcal{P}_{n}(x),~\mathcal{P}_{0}(x)=1,$ (5)
which covers more then 60 examples in OEIS (see [18, Section 4]). The exponential generating function of $\mathcal{P}_{n}(x)$ has the closed-form [18, Eq. (36)]:
$\sum_{n=0}^{\infty}\mathcal{P}_{n}(x)\frac{z^{n}}{n!}=\mathrm{e}^{p(1-x)z}\left(\frac{1-x}{1-x\mathrm{e}^{q(1-x)z}}\right)^{r}.$
Using Theorem (1) and the recursions (5) and (6), we get the following result.
Corollary 2.
Let $\mathcal{P}_{n}(x)$ be the polynomials satisfying the recursion (5). If $q+2p\geqslant qr\geqslant 2p>0$, then $\mathcal{P}_{n}(x)$ is bi-gamma-positive and $x^{n}\mathcal{P}_{n}(1/x)$ is ratio monotone. If $q(1+r)\geqslant 2p\geqslant qr>0$, then $\mathcal{P}_{n}(x)$ is ratio monotone.
This paper is organized as follows. Section 2 is devoted to the proof of Theorem 1. In Section 3, we present some applications of Theorem 1.
@@ 2.1 A proof of the bi-gamma-positivity of $f_{n}(x)$
with $\alpha_{0,0}=1,~\beta_{0,0}=0$ and $\alpha_{0,k}=\beta_{0,k}=0$ for $k\geqslant 1$. Therefore, when $a+c\geqslant b\geqslant c>0$, we see that $f_{n}(x)$ is bi-gamma-positive. Moreover, if $a>0$ and $b=c$, then $f_{n}(x)$ is reduced to a gamma-positive polynomial. ∎
Suppose that $p(x)$ and $q(x)$ have only real zeros, the zeros of $p(x)$ are $\xi_{1}\leqslant\cdots\leqslant\xi_{n}$, and that those of $q(x)$ are $\theta_{1}\leqslant\cdots\leqslant\theta_{m}$. Following [15], we say that $p(x)$ interlaces $q(x)$ if $\deg q(x)=1+\deg p(x)$ and the zeros of $p(x)$ and $q(x)$ satisfy $\theta_{1}\leqslant\xi_{1}\leqslant\theta_{2}\leqslant\cdots\leqslant\xi_{n}\leqslant\theta_{n+1}$. We use the notation $p(x)\prec q(x)$ for $p(x)$ interlaces $q(x)$. By [19, Theorem 2.1], one can immediately get the following result.
@@
Let $\{f_{n}(x)\}_{n\geqslant 0}$ be a sequence of polynomials satisfying (4). Then $f_{n}(x)$ has only real zeros and so it is unimodal.
@@ 2.2 A proof of the ratio monotonicity of $f_{n}(x)$
Comparing with (9) and (10), we get the desired result. This completes the proof. ∎
@@ 3.1 $q$-Eulerian polynomials, $1/k$-Eulerian polynomials and generalized Eulerian polynomials
Setting $L_{n}(x,q)=x^{n}A_{n+1}(1/x,q)$, it follows from (6) and (18) that
$L_{n+1}(x,q)=(nx+qx+1)L_{n}(x,q)+x(1-x)\frac{\mathrm{d}}{\mathrm{d}x}L_{n}(x,q).$ (19)
Following Hwang-Chern-Duh [18, p. 26], the polynomials $L_{n}(x,q)$ can be called LI Shanlan polynomials, since these polynomials first appeared in his 1867 book. Combining the recursions (18), (19) and Theorem 1, we can give the following result.
@@
It should be noted that $A_{n}(x,2)$ is the big descent polynomials over $\mathfrak{S}_{n+1}$, where a big descent is an index $i$ such that $\pi(i)\geqslant\pi(i+1)+2$, see [28, A120434]. We list the first few polynomials:
@@
The bi-gamma-positivity of $A_{n}^{(k)}(x)$ was first established in [22, Section 3.4], and Yan-Yang-Lin [31] gave a nice combinatorial proof of this result. Combining Corollary 5 and (20), we get the following result.
@@
These polynomials were introduced by Morisita [24], and they were also independently studied by Carlitz-Scoville [9]. It should be noted that these polynomials appeared in the context of random staircase tableaux, and Hitczenko-Janson [17] investigated their asymptotic distribution. Combining (21) and Theorem 1, we end this subsection by giving the following result.
@@ 3.3 The $r$-colored Eulerian polynomials
When $r=1$ and $r=2$, the polynomial $A_{n,r}(x)$ reduces to the types $A$ and $B$ Eulerian polynomials $A_{n}(x)$ and $B_{n}(x)$, respectively. Very recently, there has been much work devoted to the bi-gamma-positivity of the $r$-colored Eulerian polynomials and their variations, see [2, 16, 22] for details. In particular, it is now well known that $A_{n,r}(x)$ is bi-gamma-positive when $r>2$, see [2, Eq. (21)] and [22, Theorem 7.5].